Session: Jun 2013 (session 2)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Noto (C. Nocita at Wed May 29 14:10:54 2013) Minor failures:RECE
Fault of Right channel LNA. Only Left polarization working.

Medicina (A. Cattani at Wed May 29 14:15:09 2013) Success:

Yebes (Rubén Fernández at Wed May 29 15:10:17 2013) Success:
No known problems

Noto (S. Buttaccio at Thu May 30 07:17:35 2013) Failure:
Wrong frequencies setup for ND

Westerbork (Geert Kuper at Thu May 30 19:58:46 2013) Success:

Onsala (Michael Lindqvist at Fri May 31 04:51:18 2013) Success: RECO
Mark5B+ problems first scans.

Effelsberg (Georgi at Fri May 31 07:19:24 2013) Success:

Torun (K. Borkowski at Mon Jun 03 07:47:06 2013) Minor failures:
FTP fringes uncovered swapped polarization. This was rectified by cable swap before the last ftp.